# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v6

@@ IV-C Variants of Image Harmonization Task
Blind image harmonization: Most image harmonization methods require the foreground mask as input, which means that the inharmonious region is known in advance. However, in real-world applications, we may not know the exact inharmonious region in advance. Image harmonization without foreground mask is called blind image harmonization. Cun and Pun [25] considered the problem of blind image harmonization. They proposed to predict the inharmonious region mask in the attention block, which deals with the foreground and background separately according to the predicted mask. Subsequently, some works [93, 94, 193, 192] focus on inharmonious region localization task, which aims to localize the inharmonious region in an image. Liang et al. [93] explored aggregating multi-scale contextual information and suppressing redundant information. The methods [94, 193] proposed to magnify the domain discrepancy between foreground and background using color mapping for ease of identifying the inharmonious region.
@@ IV-E Datasets and Evaluation Metrics
For quantitative evaluation, existing works adopt metrics including Mean Square Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity index (SSIM) [141], Learned Perceptual Image Patch Similarity (LPIPS) [223] to calculate the distance between harmonized result and ground-truth. These metrics can also be calculated only within the foreground region. For qualitative evaluation, they conduct user study on real composite images by asking engaged users to select the most realistic images and calculate the metric (e.g., B-T score [6]).
@@ V Shadow Generation
In the previous section, image harmonization methods could adjust the foreground appearance to make it compatible with the background, but they ignore the fact that the inserted object may also have impact on the background (e.g., reflection, shadow). For example, if background objects cast shadows on the ground but the inserted object does not have shadow, the composite image would look unrealistic. To address this issue, shadow generation task aims to generate plausible shadow for the foreground object according to background illumination information to make the composite image more realistic. Similar to Section IV, we divide the existing methods into rendering based methods and non-rendering based methods.
@@ V-B Non-rendering based Methods
Other methods [225, 62, 101, 56] utilized paired training data (paired images with and without foreground shadow) to generate better shadow images. ShadowGAN [225] employed standard conditional GAN with reconstruction loss, local adversarial loss, and global adversarial loss to generate shadow for the inserted 3D foreground objects. Inoue et al. [62] developed a multi-task framework with two decoders accounting for depth map prediction and ambient occlusion map prediction respectively. ARShadowGAN [101] proposed an attention-guided residual network. The network predicts two attention maps for background shadow and occluder respectively, which are concatenated with composite image and foreground object mask to produce a residual shadow image. SGRNet [56] designed a two-stage shadow generation network. In the first stage, foreground features and background features are interacted using cross-attention to predict a shadow mask. In the second stage, they predict shadow parameters which are used to darken the input composite image. Then, the darkened image is combined with the input composite image with shadow matte. Meng et al. [116] adopted a similar two-stage pipeline and proposed to generate the shadow region by fusing multiple underexposure images. DMASNet [164] decomposed shadow mask prediction into box prediction and shape prediction, followed by attending relevant background shadow pixels to fill in the predicted shadow region. SGDiffusion [105] is the first work on shadow generation using diffusion model, which is built upon ControlNet [222] with extra intensity module to refine the shadow intensity. [189, 165] also train conditional diffusion model for shadow generation. [231, 208] first predict coarse shadow mask and then feed the shadow mask to diffusion model.
@@ V-C Datasets and Evaluation Metrics
Instead of constructing synthetic datasets [56, 105] by removing the shadows, [189] constructed real-world dataset by taking photos with object (factual image) or without object (counterfactual image). However, this approach to construct dataset is very costly and labor-intensive.
To evaluate the quality of generated composite images with foreground shadows, existing shadow generation works [211] without paired data adopt Frechet Inception Distance (FID) [54] and Manipulation Score (MS) [13] to measure the realism of generated shadow images. For the works [101, 56] with paired data, they adopt Structural SIMilarity index (SSIM) [141] and Root Mean Square Error (RMSE) [4] to measure the difference between generated image and ground-truth image. SSIM and RMSE can also be calculated only within the ground-truth foreground shadow region. Liu et al. [101] also use Balanced Error Rate (BER) [119] to evaluate the quality of predicted shadow mask based on ground-truth shadow mask.
@@ V-D Experiments
VI Generative Composition
Figure 16: The visualization results of PBE [202], ObjectStitch [152], and ControlCom [213] for generative image composition.
@@ VI Generative Composition
VI-A Deep Learning Methods
@@
The first group of methods [50, 110, 183] utilizes off-the-shelf foundation generation model, which does not require training or finetuning. They aim to generate high-quality composite images by manipulating the foreground and background elements (e.g., feature, attention) through the denoising process.
The second group of methods [202, 152, 227, 213, 209, 19, 88, 109, 163, 228, 189] learn object-to-object mapping conditioned on the background information. They train a diffusion model on abundant pairs of foregrounds and backgrounds, so that it can be directly applied to a new pair of foreground and background at test time. In the testing period, if a few images containing the foreground object are available, the pretrained model can also be finetuned on these images for better performance.
@@
In the pioneering works [202, 152], they construct massive training triplets of foregrounds, backgrounds, and ground-truth real images based on large-scale image datasets [75], in which the foregrounds are cropped from real images followed by color and geometry perturbation. Then, they adapt conditional diffusion model to this task. In particular, the background image, bounding box mask, and noisy image are concatenated as input, while the foreground is injected into the network via cross-attention. Kulal et al. [74] adopted a similar approach, but focused on human generation. Some subsequent methods focuses on enhancing the ability of detail preservation. For example, Zhang et al. [213] proposed global-and-local fusion, in which shallow foreground features are used to enhance the details. Chen et al. [19] extracted high-frequency information for better detail preservation. Different works [213, 209] also attempted to control image composition from different perspectives. For example, [209] provided the target camera viewpoint of foreground object. [213] can selectively adjust the illumination and pose of foreground object to match the background. Some methods [53, 86, 230] aimed to insert the object into any reasonable place in the background image without the provided bounding box. Some methods [230, 190] explored generating plausible shadow and reflection for the inserted foreground without the spatial constraint of bounding box.
VI-B Datasets and Evaluation Metrics
@@
Training diffusion model requires massive training triplets of foregrounds, backgrounds, and ground-truth real images. Previous works [202, 152] proposed to crop the foregrounds from real images and perturb the foregrounds (e.g., color transfer, geometric transformation), so that we can have perturbed foreground, masked background, and ground-truth real image. The multi-view datasets and video datasets can also be used to simulate more diverse and realistic geometry perturbation.
@@
In real-world application scenarios, there exist no ground-truth images for a pair of foreground and background, so we cannot calculate the distance between generated image and ground-truth image. Therefore, previous works [202, 152, 227] used FID [54] to measure the discrepancy between generated images and real images, quality score [45] to evaluate the authenticity of each image, and CLIP score [135] to measure the similarity between generated foreground and reference foreground.
VI-C Experiments
We evaluate different methods [202, 152, 213] on COCOEE test set provided by [202] and show the visualization results in Fig. 16. For [213], we show the results of its four versions: (0,0) adjusts neither illumination or pose, functioning as image blending; (1,0) only adjusts illumination, functioning as image harmonization; (0,1) only adjusts pose; (1,1) adjusts both illumination and pose like [202, 152].
We can observe that they have the potential to produce high-quality composite images, in which the foregrounds are realistic and compatible with background. Even for the drastic viewpoint change (e.g., row 2), they are capable of producing impressive results. However, in some challenging cases, the generated results may lose the details of foreground object (e.g., row 3, row 4) or contain noticeable artifacts (e.g., row 6). Compared with [202, 152], [213] is more adept at preserving foreground details, but is also more likely to produce artifacts and unrealistic objects. It is still very challenging to achieve both dramatic viewpoint/pose variation and foreground detail preservation.
@@ VI-C Experiments
Additionally, these methods require manual specification of reasonable bounding box and cannot generate shadow that falls out of the bounding box, so object placement and shadow generation are actually not fully accomplished.
VII Foreground Object Search
@@ VII-A Traditional Methods
VII-B Deep Learning Methods
@@
VII-C Datasets and Evaluation Metrics

### v7

@@ IV-C Variants of Image Harmonization Task
Blind image harmonization: Most image harmonization methods require the foreground mask as input, which means that the inharmonious region is known in advance. However, in real-world applications, we may not know the exact inharmonious region in advance. Image harmonization without foreground mask is called blind image harmonization. Cun and Pun [29] considered the problem of blind image harmonization. They proposed to predict the inharmonious region mask in the attention block, which deals with the foreground and background separately according to the predicted mask.
Subsequently, some works [103, 104, 212, 211, 247, 161] focused on inharmonious region localization task, which aims to localize the inharmonious region in an image. Liang et al. [103] explored aggregating multi-scale contextual information and suppressing redundant information. The methods [104, 212] proposed to magnify the domain discrepancy between foreground and background using color mapping for ease of identifying the inharmonious region.
@@ IV-E Datasets and Evaluation Metrics
Existing works adopt metrics including Mean Square Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity index (SSIM) [154], Learned Perceptual Image Patch Similarity (LPIPS) [245] to calculate the distance between harmonized result and ground-truth. These metrics can also be calculated only within the foreground region. Besides, they conduct user study on real composite images by asking engaged users to select the most realistic images and calculate the metric (e.g., B-T score [7], ratio).
@@ IV-F Experiments
For practical usage, in usual lighting conditions, color-to-color transformation methods like [52] are recommended, because they can process high-resolution images efficiently and preserve the image details well. In unusual lighting conditions (e.g., neon light), diffusion-based methods may perform better.
@@
For practical usage, in most cases, feed-forward GAN-based methods like [11] can achieve satisfactory results. In the challenging cases such as dense textures or abstract style, feed-forward diffusion-based methods like [119] are able to demonstrate their strengths.
@@ V Shadow Generation
In the previous section, image harmonization methods could adjust the foreground appearance to make it compatible with the background, but they ignore the fact that the inserted object may also have impact on the background. For example, if background objects cast shadows on the ground but the inserted object does not have shadow, the composite image would look unrealistic. To address this issue, shadow generation task aims to generate plausible shadow for the foreground object according to background illumination information to make the composite image more realistic. Similar to Section IV, we divide the existing methods into rendering based methods and non-rendering based methods.
@@ V-B Non-rendering based Methods
Other methods [248, 69, 111, 62] utilized paired training data (paired images with and without foreground shadow) to generate better shadow images. ShadowGAN [248] employed standard conditional GAN with reconstruction loss, local adversarial loss, and global adversarial loss to generate shadow for the inserted 3D foreground objects. Inoue et al. [69] developed a multi-task framework with two decoders accounting for depth map prediction and ambient occlusion map prediction respectively. ARShadowGAN [111] proposed an attention-guided residual network. The network predicts two attention maps for background shadow and occluder respectively, which are concatenated with composite image and foreground object mask to produce a residual shadow image. SGRNet [62] designed a two-stage shadow generation network. In the first stage, foreground features and background features are interacted using cross-attention to predict a shadow mask. In the second stage, they predict shadow parameters which are used to darken the input composite image. Then, the darkened image is combined with the input composite image with shadow matte. Meng et al. [127] adopted a similar two-stage pipeline and proposed to generate the shadow region by fusing multiple underexposure images. DMASNet [179] decomposed shadow mask prediction into box prediction and shape prediction, followed by attending relevant background shadow pixels to fill in the predicted shadow region. SGDiffusion [115] is the first work on shadow generation using diffusion model, which is built upon ControlNet [243] with extra intensity module to refine the shadow intensity. [208, 180, 196] also trained conditional diffusion model for shadow generation. [256, 229] first predicted coarse shadow mask and then fed the shadow mask to diffusion model. Zhao et al. [252] injected geometry prior, i.e., shadow location and shape, into diffusion model to enhance the quality of generated shadows with complex shapes.
@@ V-C Datasets and Evaluation Metrics
Instead of constructing synthetic datasets [62, 115] by removing the shadows, [208, 79] constructed real-world dataset by taking photos with object (factual image) or without object (counterfactual image). However, this approach to construct dataset is very costly and labor-intensive.
To evaluate the quality of generated composite images with foreground shadows, 1) existing shadow generation works [232] without paired data adopt Frechet Inception Distance (FID) [60] and Manipulation Score (MS) [15] to measure the realism of generated shadow images. 2) For the works [111, 62] with paired data, they adopt Structural SIMilarity index (SSIM) [154] and Root Mean Square Error (RMSE) [5] to measure the difference between generated image and ground-truth image. SSIM and RMSE can also be calculated only within the ground-truth foreground shadow region. Liu et al. [111] also use Balanced Error Rate (BER) [130] to evaluate the quality of predicted shadow mask based on ground-truth shadow mask. 3) User study is also called for to ensure that generated shadows comply with human perception. Participants are asked to select the most realistic images, based on which some metrics (e.g., B-T score [7], ratio) are calculated.
@@ V-D Experiments
For practical usage, foundation diffusion model (e.g, SD [152], FLUX [1]) is imperative to generate realistic shadows. More advanced foundation model usually lead to better performance, especially in the challenging cases.
VI Reflection Generation
Figure 16: We show four examples from DEROBA dataset. From left to right in each example, we show the composite image without foreground reflection, the foreground mask, and the ground-truth image with foreground reflection.
Besides shadow, reflection is another impact that the foreground object may have on the background, especially when the foreground object is above the water or on the desktop with reflective material. To address this issue, reflection generation task aims to generate plausible reflection for the foreground object to make the composite image more realistic. Similar to Section V, we divide the existing methods into rendering based methods and non-rendering based methods. ’
VI-A Rendering based Methods
Ma et al. [125] first estimated the lighting information and then employed renderer to generate coarse reflection, which is further refined using neural network.
VI-B Non-rendering based Methods
Some works treat reflection generation as image-to-image translation task without estimating lighting/geometry information or employing renderer. For example, the works [208, 180] employed conditional diffusion model to synthesize reflection for the inserted foreground object. The works [31, 32] adopt similar approaches but focused on mirror reflection.
VI-C Datasets and Evaluation Metrics
Training conditional diffusion model for reflection synthesis requires paired training data, i.e., a composite image without foreground reflection and a ground-truth image with foreground reflection.
Some works [31, 32] constructed rendered datasets with paired data by inserting a virtual object into 3D scene and generating reflection in the mirror with rendering technique. For example, Dhiman et al. [31] released SynMirror dataset with 66,068 objects and 198,204 rendered images.
Different from rendering pipeline and similar to the pipeline of constructing shadow generation dataset DESOBAv2 [115], DEROBA manually annotated object-reflection masks and employed the off-the-shelf inpainting model to inpaint the reflection regions, resulting in synthesized composite images.
Instead of constructing synthetic datasets [31, 32], [208, 79] constructed real-world dataset by taking photos with object (factual image) or without object (counterfactual image). However, this approach to construct dataset is very costly and labor-intensive.
To evaluate the quality of generated composite images with foreground reflections, Mean Square Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity index (SSIM) [154], Learned Perceptual Image Patch Similarity (LPIPS) [245] can be adopted to measure the difference between generated image and ground-truth image. User study is also necessary to ensure that generated reflections conform to human perception. Participants are asked to select the most realistic images, based on which some metrics (e.g., B-T score [7], ratio) are calculated.
VI-D Experiments
We compare different methods on DEROBA dataset. The baseline is ControlNet [243] taking composite image and foreground mask as the input of control encoder. Besides, we extend ControlNet to RGDiffusion considering the property of reflection generation task. Specifically, we inject the cropped foreground into denoising UNet via cross-attention, considering that reflections are usually the horizontal or vertical mirror images of foreground objects. We show the results of ControlNet and RGDiffusion in Fig. 17, based on which RGDiffusion has more accurate shapes and clearer details.
Figure 17: The visualization results of different reflection generation methods on DEROBA dataset. From left to right in each group, we show the input composite image, the composite foreground mask, the generated results of ControlNet [243], RGDiffusion, and the ground-truth reflection image.
VII Generative Composition
Figure 18: The visualization results of different generative composition methods on MureCom [120] dataset. From left to right in each row, we show the background with foreground bounding box, five reference images of the same foreground object, the generated results of PbE [222], Anydoor [22], ControlCom [234], MureObjectStitch [20], and InsertAnything [166].
@@ VII Generative Composition
VII-A Deep Learning Methods
@@ VII-A1 Training-free Methods
The first group of methods [56, 121, 201, 96, 219] utilizes off-the-shelf foundation generation model, which does not require training or finetuning. They aim to generate high-quality composite images by manipulating the foreground and background elements (e.g., feature, attention) through the denoising process.
VII-A2 Training-based Methods
The second group of methods [222, 167, 250, 234, 230, 22, 98, 120, 178, 251, 208] require training or finetuning. They can be further divided into two subgroups, according to whether object-specific finetuning is indispensable. The first subgroup of methods [222, 167, 250, 234, 230, 22] train a diffusion model on abundant pairs of foregrounds and backgrounds, so that it can be directly applied to a new pair of foreground and background at test time. Their models need to take one or more reference images of the inserted foreground object as input. In the testing period, if a few images containing the foreground object are available, we can optionally finetune the pretrained model on these images and may achieve better performance. The second subgroup of methods [23, 120, 98, 153] do not train the model on large-scale dataset. Most of them associate the target object with one rare token, so object-specific finetuning (training on a few images of the same foreground object) is indispensable for these methods. Because the first subgroup is dominant, we will mainly introduce the first subgroup of methods.
@@ VII-A2 Training-based Methods
Among the first subgroup of methods, the pioneering works like PbE [222] and ObjectStitch [167] construct massive training triplets of foregrounds, backgrounds, and ground-truth real images based on large-scale image datasets [83], in which the foregrounds are cropped from real images followed by color and geometry perturbation. Then, they adapt conditional diffusion model to this task. In particular, the background image, bounding box mask, and noisy image are concatenated as input, while the foreground is injected into the network via cross-attention. ObjectStitch [167] is extended to MureObjectStitch [20] to support multiple foreground reference images. Kulal et al. [82] adopted a similar approach, but focused on human generation. Some subsequent methods focuses on enhancing the ability of detail preservation. For example, Zhang et al. [234] proposed global-and-local fusion, in which shallow foreground features are used to enhance the details. Chen et al. [22] extracted high-frequency information for better detail preservation. Yu et al. [228] designed a cycle formed by adding object and removing object, which is expected to facilitate each other. Recently, inspired by in-context learning, several recent works [191, 166, 67] explored image composition task under the framework of in-context learning. Specifically, noisy latent, background, foreground are equally fed into diffusion model and interact with each other. Different types of input are added with different levels of noise.
Different works [234, 230] also attempted to control image composition from different perspectives. For example, [230] provided the target camera viewpoint of foreground object. [234] can selectively adjust the illumination and pose of foreground object to match the background. Some methods [59, 95, 254] aimed to insert the object into any reasonable place in the background image without the provided bounding box. Some methods [254, 209] explored generating plausible shadow and reflection for the inserted foreground without the spatial constraint of bounding box.
VII-B Datasets and Evaluation Metrics
@@
Training diffusion model requires massive training triplets of foregrounds, backgrounds, and ground-truth real images. Previous works [222, 167] proposed to crop the foregrounds from real images and perturb the foregrounds (e.g., color transfer, geometric transformation), so that we can have perturbed foreground, masked background, and ground-truth real image. The multi-view datasets and video datasets can also be used to simulate more diverse and realistic geometry perturbation. In particular, they replace the perturbed foreground with the foreground image from another view or another video frame.
Another way to construct image composition dataset is capturing two images with or without the foreground object for the same scene [209, 79]. Specifically, they first capture an image for the background scene. Then, they place a foreground object and capture an image with the same camera viewpoint again. However, the scale of such datasets is limited by the high cost of image collection.
@@
In real-world application scenarios, there exist no ground-truth images for a pair of foreground and background, so we cannot calculate the distance between generated image and ground-truth image. Therefore, previous works [222, 167, 250] used FID [60] to measure the distribution discrepancy between generated images and real images. Quality scores [51, 205] are used to evaluate the authenticity of each image. CLIP score [147] or DINO score [13] are used to measure the similarity between generated foreground and reference foreground. User study should also be conducted to evaluate different aspects for the generated images.
VII-C Experiments
We evaluate different methods [222, 22, 234, 20, 166] on MureCom [120] dataset and show the visualization results in Fig. 18. MureCom [120] dataset has five reference images for each foreground object. For [222, 22, 234, 166] which only supports one foreground reference image, we use foreground object search methods like [236] (see Section VIII) to find the most suitable foreground reference image given the background. For [20] which supports multiple foreground reference images, we use all five foreground reference images. For [222, 22, 234, 166], we directly use their released models (“harmonization” mode for [234]). For [20], we finetune the pretrained model on each foreground object for better performance.
From Fig. 18, we can see that [222] can generate foreground object with reasonable view/pose but fails to preserve the appearance details. Compared with [22, 234, 20], [166] can preserve very small details (e.g., the letters on the cake in row 3, the pattern on the waste container in row 5), probably due to the detail preservation ability of in-context learning. The methods [22, 234, 166] exhibit notable copy-and-paste effect and heavily rely on suitable foreground reference image. If the view/pose of foreground reference image is not compatible with the background, these methods [22, 234, 166] are weak in adjusting the foreground view/pose to match the background.
@@ VII-C Experiments
Additionally, these methods require manual specification of reasonable bounding box and most of them cannot generate high-quality shadow/reflection that falls out of the bounding box, so object placement and shadow/reflection generation are actually not fully accomplished.
VIII Foreground Object Search
@@ VIII-A Traditional Methods
VIII-B Deep Learning Methods
@@
VIII-C Datasets and Evaluation Metrics